n 2: Suppose a car speedometer has a 4.4 % uncertainty.

a. what is the car's highest possible speed, in kilometers per hour, when the speedometer reads 94km/h?

b. what is the car's lowest possible speed, in kilometers per hour, when the speedometer reads 94km/h?
Answer #1

Percentage uncertainty formula give as

% = [ (delta v)/v] * 100

Delta v = 4.4*94/100 = 4.13 km/h

a) highest possible speed = 94 + 4.13 = 98.13 km/h

b) lowest possible speed = 94 - 4.13 = 89.87 Km/h